OpenStudy (anonymous):

(_____)^2 = (csc x-1)(csc x+1) fill in the blank

OpenStudy (ikram002p):

Hint :- (a-b)(a+b)=a^2-b^2

OpenStudy (ikram002p):

so \(\large (\csc x-1)(\csc x+1) =\csc ^2 -1^2=\csc ^2 -1\)

OpenStudy (ikram002p):

nw are you familiar with this ? \(\large \cot^2(x) + 1 = \csc^2(x)\) so \(\large \cot^2(x) = \csc^2(x)-1\) got it ?

OpenStudy (ikram002p):

let me know if u dint understand it yet :D

OpenStudy (anonymous):

so its cot^2(x)?

OpenStudy (ikram002p):

yep :D

OpenStudy (anonymous):

thanks!

OpenStudy (ikram002p):

np :)